# SLS-BRD: A system-level approach to seeking generalised feedback Nash equilibria

### v2

@@ III-A System-level best-response mappings
The sets $(\mathcal{C}_{x,n},\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ ($\forall p\in\mathcal{P}$) are designed to impose some structure directly on the policy $\bm{K}^{p}$ (Corollary 4.1), often in the form of sparsity constraints. An important class of such structural constraints encodes information patterns incurred by the presence of actuation and sensing delays: Let $\bm{K}^{p}\in\bm{\mathcal{C}}^{p}$ ($\forall p$) satisfy the restrictions
$\bm{\mathcal{C}}^{p}=\{\bm{K}^{p}\in\mathcal{L}(\bm{\mathcal{X}},\bm{\mathcal{U}}^{p}):$ Actions [$\text{Actions }[B^{p}(K^{p}x)_{t}]_{i}\text{ respectively}$ (39)
affect and feedback the component
$[x_{t}]_{i}\text{ with the delays }d_{a},d_{s}>0\},$
and consider the operators $\bm{S}_{x}:\bm{\Phi_{x}}\mapsto(S_{x,n}\odot\Phi_{x,n})_{n\in\mathbb{N}_{+}}$ and $\bm{S}_{u}^{p}:\bm{\Phi_{u}}^{p}\mapsto(S_{u,n}^{p}\odot\Phi_{u,n}^{p})_{n\in\mathbb{N}_{+}}$ ($\forall p$), given the matrices
$(S_{x,n})_{n\in\mathbb{N}_{+}}$ $=\left(\texttt{Sp}(A^{\max{(0,\lfloor\frac{n-d_{a}}{d_{s}}\rfloor)}})\right)_{n\in\mathbb{N}_{+}},$
$(S_{u,n}^{p})_{n\in\mathbb{N}_{+}}$ $=\left(\texttt{Sp}({B^{p}}^{{\mkern-1.5mu\mathsf{T}}}A^{\max{(0,\lfloor\frac{n-d_{a}}{d_{s}}\rfloor)}})\right)_{n\in\mathbb{N}_{+}},$
with $\texttt{Sp}(\cdot)$ denoting the sparsity pattern of a matrix [^1] The operator $\texttt{Sp}(A)$ produces a binary matrix such that $[\texttt{Sp}(A)]_{i,j}=1$ if $[A]_{i,j}\neq 0$ and $[\texttt{Sp}(A)]_{i,j}=0$ otherwise.. It can be shown that $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}\in\bm{\mathcal{C}}^{p}$ ($\forall p$) if its parametrisation satisfy $(\Phi_{x,n}\in\mathcal{C}_{x,n})_{n\in\mathbb{N}^{+}}$ and $(\Phi^{p}_{u,n}\in\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ with
@@
The constraints Eq. (40) enforce that the closed-loop response to the noise obeys an information pattern induced by the dynamics of the game $\mathcal{G}_{\infty}^{\text{LQ}}$. Specifically, $[S_{x,n}]_{i,j}=0$ (respectively, $[S_{u,n}^{p}]_{i,j}=0$) indicates that any disturbance to the $j$-th component of the state, $[x_{t}]_{j}$, does not affect $[x_{t+n}]_{i}$ (does not feedback into action $[u_{t+n}^{p}]_{i}$). Finally, note that $(\mathcal{C}_{x,n},\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ are compact convex sets, as they correspond to the kernels of the linear operators ($I-\bm{S}_{x})$ and $(I-\bm{S}_{u}^{p}$).
The ability to impose a desired policy structure using convex constraints is a central feature of the SLS framework. In the context of dynamic games, it allows for describing and, most importantly, solving problems where players have asymmetric information patterns; a major challenge for feedback Nash equilibrium problems [43, 41]. For $(\mathcal{C}_{x,n},\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ from Eq. (40), conditions for the existence of GFNE (i.e., $\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}\neq\emptyset$), can be stated in terms of the delay parameters $d_{a}$ and $d_{s}$. Unless stated otherwise, the mappings $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ (and $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$) are assumed to include these constraints with $d_{a}=d_{s}=1$
III-B System-level best-response dynamics
@@
A policy profile $\bm{K}^{*}=(\bm{\Phi_{u}}^{1^{*}},\ldots,\bm{\Phi_{u}}^{N_{P}^{*}})\bm{\Phi_{x}}^{*^{-1}}\in\bm{\mathcal{C}}$, is a GFNE of $\mathcal{G}_{\infty}^{\text{LQ}}$ if $\bm{\Phi_{u}}^{*}\in BR_{\Phi}(\bm{\Phi_{u}}^{*})$, or, equivalently,
$\bm{\Phi_{u}}^{p^{*}}\in BR_{\Phi}^{p}(\bm{\Phi_{u}}^{-p^{*}}),\quad\forall p\in\mathcal{P}.$ (41)
@@
Consider an arbitrary fixed-point $\bm{\Phi_{u}}^{*}\in BR_{\Phi}^{p}(\bm{\Phi_{u}}^{*})$. From Theorem 4, we have $\bm{\Phi_{x}}^{*}=\bm{F_{\Phi}}\bm{\Phi_{u}}^{*}$. Now, consider policies $\bm{K}^{p^{*}}=\bm{\Phi_{u}}^{p^{*}}(\bm{\Phi_{x}}^{*})^{-1}$, $p\in\mathcal{P}$. Clearly, $\bm{\Phi_{u}}^{p^{*}}=\bm{K}^{p^{*}}\bm{\Phi_{x}}^{*}$. As a consequence, for any $\bm{w}\in\bm{\mathcal{W}}$,
$\bm{\Phi_{u}}^{p^{*}}\bm{w}=\bm{K}^{p^{*}}\bm{\Phi_{x}}^{*}\bm{w}\leavevmode\nobreak\ \leavevmode\nobreak\ \Longleftrightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ \bm{u}^{p^{*}}=\bm{K}^{p^{*}}\bm{x}^{*}.$
and, by definition, $\bm{\Phi_{u}}^{p^{*}}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p^{*}})$ imply $\bm{u}^{p^{*}}\in U^{p}(\bm{u}^{-p^{*}})$. Thus, $\bm{u}^{*}$ is the open-loop realisation of the policy $\bm{K}^{*}$. Finally, since $J^{p}(\bm{u}^{p^{*}},\bm{u}^{-p^{*}})\cong J^{p}(\bm{\Phi_{u}}^{p^{*}},\bm{\Phi_{u}}^{-p^{*}})$ and $\bm{\Phi_{u}}^{*}\in\Omega_{\mathcal{G}_{\infty}^{\Phi}}$, we conclude that no player can obtain an admissible policy that unilaterally improves its cost, that is, $\bm{K}^{*}\in\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}$. ∎
The relationship between $BR$ and $BR_{\Phi}$ implies that a GFNE of $\mathcal{G}_{\infty}^{\text{LQ}}$ can be obtained analytically from a GNE of $\mathcal{G}_{\infty}^{\Phi}$. This fact allows us to adapt the BRD-GFNE procedure (Algorithm 3) and propose a procedure for GFNE seeking in constrained infinite-horizon dynamic games based on the mappings $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$. This system-level best-response dynamics (SLS-BRD) approach is given in Algorithm 4. We remark on some technical aspects:
  • The pair $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{u,k}}^{p})$ defines the $p$-th player’s parametrisation after $k\in\mathbb{N}$ updates, that is, $\bm{\Phi_{x,k}}^{p}=(\Phi_{x,k,t}^{p})_{t\in\mathbb{N}}$ and $\bm{\Phi_{u,k}}^{p}=(\Phi_{u,k,t}^{p})_{t\in\mathbb{N}}$. The update index $k\in\mathbb{N}$ should not be mistaken for the stage index $t\in\mathbb{N}$.
  • Updating the policy $\bm{K}_{k+1}^{p}$ consists of either employing Corollary 4.1 with maps $(\bm{\Phi_{x,k+1}}^{p},\bm{\Phi_{u,k+1}}^{p})$, or directly computing its kernel $\bm{\Phi_{k+1}}^{p}=([M_{\Phi_{u,k+1}^{p}}]_{n}\Phi_{x,k+1}^{-1})_{n\in\mathbb{N}}$.
@@
Output: GFNE $\bm{K}^{*}=(\bm{K}^{1^{*}},\ldots,\bm{K}^{N_{P}^{*}})$
@@
2 Initialize $\bm{K}_{0}=(\bm{K}_{0}^{1},\ldots,\bm{K}_{0}^{N_{P}})$ and $k=0$;
@@ 8            $k\coloneqq k+1$;
The SLS-BRD induces an operator $T_{\Phi}=(1-\eta)I+\eta BR_{\Phi}$, which defines the global update $\bm{\Phi_{u,k+1}}$ from $\bm{\Phi_{u,k}}$. As before, $T_{\Phi}$ and $BR_{\Phi}$ share the same fixed-points: The GNEs $\Omega_{\mathcal{G}_{\infty}^{\Phi}}$. From Theorem 6, convergence to a response $\bm{\Phi_{u}}^{*}\in T_{\Phi}(\bm{\Phi_{u}}^{*})$ then implies convergence to a policy $\bm{K}^{*}\in\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$. Hence, this learning dynamics is a formal procedure for GFNE seeking.
@@
The convergence of Algorithm 4 depends on $BR_{\Phi}$ (or $\widehat{BR}_{\Phi}$, for its tractable version) being at least non-expansive (Theorems 23). Formally,
Theorem 7.
Let the map $\widehat{BR}_{\Phi}:\bm{\mathcal{C}}_{u}\rightrightarrows\bm{\mathcal{C}}_{u}$ be $L_{\widehat{BR}_{\Phi}}$-Lipschitz, with $L_{\widehat{BR}_{\Phi}}<1$. Then, the SLS-BRD $\bm{\Phi_{u,k+1}}=\widehat{T}_{\Phi}(\bm{\Phi_{u,k}})$ converge to the unique $\varepsilon$-GNE $\bm{\Phi_{u}}^{*}\in\Omega_{\mathcal{G}_{\infty}^{\Phi}}^{\varepsilon}$ with rate
$\cfrac{\|\bm{\Phi_{u,k}}-\bm{\Phi_{u}}^{*}\|_{\ell_{2}}}{\|\bm{\Phi_{u,0}}-\bm{\Phi_{u}}^{*}\|_{\ell_{2}}}\leq\big{(}(1{-}\eta)-\eta L_{\widehat{BR}_{\Phi}}\big{)}^{k}$ (42)
@@
$\underset{\bm{\Phi_{u}}^{p}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})}{\text{minimize}}\leavevmode\nobreak$ $\mathrm{Tr}\big{[}{\bm{\Phi_{u}}^{p}}^{{\mkern-1.5mu\mathsf{T}}}\bm{H}^{pp}\bm{\Phi_{u}}^{p}$
$\hskip 40.00006pt+2(\textstyle\sum_{\tilde{p}\in\mathcal{P}\backslash\{p\}}\bm{H}^{p\tilde{p}}\bm{\Phi_{u}}^{\tilde{p}}+\bm{H}^{p0})^{{\mkern-1.5mu\mathsf{T}}}\bm{\Phi_{u}}^{p}\big{]}$ (44)
The maps $BR^{p}(\bm{\Phi_{u}}^{-p})$, $p\in\mathcal{P}$, thus correspond to the solution of quadratic programs with convex constraints $U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})$, for $\bm{\Phi_{u}}^{-p}\in\bm{\mathcal{C}}_{u}^{-p}$. In the case of $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|w_{t}\|_{\infty}\leq 1\}$, and without structural constraints, that is, $\bm{S}_{x}=I$ and $\bm{S}_{u}^{p}=I$,
@@
$\Phi_{x,{n+1}}=A\Phi_{x,n}{+}\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}\Phi^{\tilde{p}}_{u,n},\leavevmode\nobreak\ \Phi_{x,1}=I_{N_{x}},$
$\textstyle\sum_{n=0}^{N}\big{\|}([G_{x}]_{i}\Phi_{x,n})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq[g_{x}]_{i},$
$\textstyle\sum_{n=0}^{N{-}1}\big{\|}([G_{u}^{p}]_{j}\Phi_{u,n}^{p})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq[g_{u}^{p}]_{j},$
$\textstyle\sum_{n=0}^{N{-}1}\big{\|}\textstyle\sum_{\tilde{p}\in\mathcal{P}}([G_{\mathcal{G}}^{\tilde{p}}]_{l}\Phi_{u,n}^{\tilde{p}})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq[g_{\mathcal{G}}]_{l}\}.$
Using standard techniques from optimisation, the constraints Eq. (45) can be incorporated into Problem (44) as linear inequalities. The solutions for these problems, that is, the best-responses $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, are thus piecewise affine in $\bm{\Phi_{u}}^{-p}\in\bm{\mathcal{C}}_{u}^{-p}$ [44]. Consequently, also $\widehat{BR}_{\Phi}$ must be piecewise affine. A local Lipschitz constant can then be derived for each region of $\bm{\mathcal{C}}_{u}^{-p}$ that leads to a subset of the operational constraints being active. For non-generalised games (i.e., when $\mathcal{X}=\mathbb{R}^{N_{x}}$ and $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$), the structure of this affine mapping can be exploited to derive a global Lipschitz constant for $\widehat{BR}_{\Phi}$.
@@
Consider $\mathcal{X}=\mathbb{R}^{N_{x}}$ and $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$. Then, The map $\widehat{BR}_{\Phi}$ is $L_{\widehat{BR}_{\Phi}}$-Lipschitz with
$L_{\widehat{BR}_{\Phi}}=\sqrt{\textstyle\sum_{p\in\mathcal{P}}(L_{\widehat{BR}_{\Phi}}^{p})^{2}},$ (46)
given the player-specific constants
$L_{\widehat{BR}_{\Phi}}^{p}=\big{(}1+\kappa(\bm{H}^{pp})\big{)}\|(\bm{H}^{pp})^{\dagger}\bm{H}^{p,-p}\|_{2\to 2},$ (47)
with condition number $\kappa(\bm{H}^{pp})=\|(\bm{H}^{pp})^{\dagger}\|_{2\to 2}\|\bm{H}^{pp}\|_{2\to 2}$.
@@
The proof of Theorem 8 is extensive: The reader is referred to Appendix -C for the full details. Importantly, this Lipschitz constant is not tight and thus $L_{\widehat{BR}_{\Phi}}<1$ using Eq. (46) is only a sufficient (but not necessary) condition for Theorem 7 to hold in practice. Regardless, it highlights some intuitive, but non-trivial, facts about the convergence of the SLS-BRD:
@@
  • The condition $L_{\widehat{BR}_{\Phi}}<1$ is implies that the block-operator $\bm{H}=[H^{p\tilde{p}}]_{p,\tilde{p}\in\mathcal{P}}$ is diagonally dominant. This highlights the relationship between the SLS-BRD and the class of Jacobi iterative methods, where diagonal dominance of the linear system being solved is a strict requirement.
@@
  • The convergence rate of the SLS-BRD is inversely proportional to $\|D^{pp}\|_{2}^{2}$ ($\forall p$) through $\bm{H}^{pp}$. As such, faster convergence is expected for games in which players apply strong penalties to their own actions.
  • The convergence rate of the SLS-BRD depends on the number of players: Let $L_{\widehat{BR}_{\Phi}}^{p}=\sqrt{\alpha+\beta^{p}}$ for all $p\in\mathcal{P}$ given some $\alpha,\beta^{p}>0$. Then, $L_{\widehat{BR}_{\Phi}}^{2}=(\alpha N_{p})^{2}\sum_{p}\beta^{p}$. In large-scale games, players might need to become more conservative in order to ensure that $\widehat{BR}$ is a contraction.

### v3

@@ 3.1 System-level best-response mappings
The sets $(\mathcal{C}_{x,n},\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ ($\forall p\in\mathcal{P}$) are designed to impose some structure directly on the policy $\bm{K}^{p}$ (Corollary 1.1), often in the form of sparsity constraints. We consider the class of structural constraints which encode information patterns incurred by actuation and communication delays: Let $\bm{K}^{p}\in\bm{\mathcal{C}}^{p}$ ($\forall p$) satisfy the restrictions
$\bm{\mathcal{C}}^{p}=\{\bm{K}^{p}\in\mathcal{L}(\ell_{\infty}^{N_{x}},\ell_{\infty}^{N_{u}^{p}}):$ The state [$\text{The state }[x_{t}]_{i}\text{ is propagated to}$
and affected by [$\text{and affected by }[B^{p}K^{p}x_{t}]_{i}\text{ with}$
$\text{delays }d_{c},d_{a}>0\text{, respectively}\}.$
and consider the operators $\bm{S}_{x}:\bm{\Phi_{x}}\mapsto(S_{x,n}\odot\Phi_{x,n})_{n\in\mathbb{N}_{+}}$ and $\bm{S}_{u}^{p}:\bm{\Phi_{u}}^{p}\mapsto(S_{u,n}^{p}\odot\Phi_{u,n}^{p})_{n\in\mathbb{N}_{+}}$ ($\forall p\in\mathcal{P}$), given the signals
$(S_{x,n})_{n\in\mathbb{N}_{+}}$ $=\left(\texttt{Sp}(A^{\max{(0,\lfloor\frac{n-d_{a}}{d_{c}}\rfloor)}})\right)_{n\in\mathbb{N}_{+}};$
$(S_{u,n}^{p})_{n\in\mathbb{N}_{+}}$ $=\left(\texttt{Sp}({B^{p}}^{{\mkern-1.5mu\mathsf{T}}}A^{\max{(0,\lfloor\frac{n-d_{a}}{d_{c}}\rfloor)}})\right)_{n\in\mathbb{N}_{+}},$
with $\texttt{Sp}(\cdot)$ denoting the sparsity pattern of a matrix, that is, $[\texttt{Sp}(X)]_{i,j}=1$ if $[X]_{i,j}\neq 0$ and $[\texttt{Sp}(X)]_{i,j}=0$ otherwise, for any matrix $X$. It can be shown that $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}\in\bm{\mathcal{C}}^{p}$ if $(\Phi_{x,n}\in\mathcal{C}_{x,n})_{n\in\mathbb{N}^{+}}$ and $(\Phi^{p}_{u,n}\in\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ with convex sets
@@
The constraints Eq. (38) enforce that the closed-loop response to the noise obeys an information pattern induced by the dynamics of the game $\mathcal{G}_{\infty}^{\text{LQ}}$. Specifically, $[S_{x,n}]_{i,j}=0$ (resp., $[S_{u,n}^{p}]_{i,j}=0$) implies that disturbances to the $j$-th component of the state, $[x_{t}]_{j}$, should not affect the state $[x_{t+n}]_{i}$ (the action $[u_{t+n}^{p}]_{i}$) when the policies $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}$ are employed. Enforcing sparsity also benefits the tractability of the best-response maps, as this reduces the effective number of decision variables.
The ability to impose a desired policy structure using convex constraints is a central feature of the SLS framework. In the context of dynamic games, it allows for describing and, most importantly, solving problems where players have asymmetric information patterns; a major challenge for feedback Nash equilibrium problems [44, 39]. Considering $(\mathcal{C}_{x,n},\mathcal{C}^{p}_{u,n})_{n\in\mathbb{N}^{+}}$ from Eq. (38), the feasibility of the best-response maps $BR^{p}$ ($\forall p$) become dependent on the parameters $d_{a}$ and $d_{c}$: The larger their values the more restrictive the search space of matrices $\{\Phi_{x,n},\Phi_{u,n}^{p}\}_{n=1}^{N}$. As such, the actuation and communication delays associated with the game directly affect the existence of GFNE (that is, $\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}\neq\emptyset$). In practice, whether these structural constraints are consistent or not can be assessed by solving the feasibility problem associated with Problem (30) [43]. A formal analysis of the requirements for $d_{a}$ and $d_{c}$ to ensure $\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}\neq\emptyset$ is beyond the scope of this paper.
3.2 System-level best-response dynamics
@@
A policy profile $\bm{K}^{\star}=(\bm{\Phi_{u}}^{1^{\star}},\ldots,\bm{\Phi_{u}}^{N_{P}^{\star}})\bm{\Phi_{x}}^{\star^{-1}}\in\bm{\mathcal{C}}$, is a GFNE of $\mathcal{G}_{\infty}^{\text{LQ}}$ if $\bm{\Phi_{u}}^{\star}\in BR_{\Phi}(\bm{\Phi_{u}}^{\star})$, or, equivalently,
$\bm{\Phi_{u}}^{p^{\star}}\in BR_{\Phi}^{p}(\bm{\Phi_{u}}^{-p^{\star}}),\quad\forall p\in\mathcal{P}.$ (39)
@@
Consider an arbitrary fixed-point $\bm{\Phi_{u}}^{\star}\in BR_{\Phi}^{p}(\bm{\Phi_{u}}^{\star})$. From Theorem 1, we have $\bm{\Phi_{x}}^{\star}=\bm{F_{\Phi}}\bm{\Phi_{u}}^{\star}$. Now, consider policies $\bm{K}^{p^{\star}}=\bm{\Phi_{u}}^{p^{\star}}(\bm{\Phi_{x}}^{\star})^{-1}$, $p\in\mathcal{P}$. Clearly, $\bm{\Phi_{u}}^{p^{\star}}=\bm{K}^{p^{\star}}\bm{\Phi_{x}}^{\star}$. As a consequence, for any $\bm{w}\in\bm{\mathcal{W}}$,
$\bm{\Phi_{u}}^{p^{\star}}\bm{w}=\bm{K}^{p^{\star}}\bm{\Phi_{x}}^{\star}\bm{w}~{}~{}\iff~{}~{}\bm{u}^{p^{\star}}=\bm{K}^{p^{\star}}\bm{x}^{\star}.$
and, by definition, $\bm{\Phi_{u}}^{p^{\star}}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p^{\star}})$ imply $\bm{u}^{p^{\star}}\in U^{p}(\bm{u}^{-p^{\star}})$. Thus, $\bm{u}^{\star}$ is an open-loop realisation of the policy $\bm{K}^{\star}$. Finally, since $J^{p}(\bm{u}^{p^{\star}},\bm{u}^{-p^{\star}})\cong J^{p}(\bm{\Phi_{u}}^{p^{\star}},\bm{\Phi_{u}}^{-p^{\star}})$ and $\bm{\Phi_{u}}^{\star}\in\Omega_{\mathcal{G}_{\infty}^{\Phi}}$, we conclude that no player can obtain an admissible policy that unilaterally improves its cost, that is, $\bm{K}^{\star}\in\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}$. ∎
@@
Output: GFNE $\bm{K}^{\star}=(\bm{K}^{1^{\star}},\ldots,\bm{K}^{N_{P}^{\star}})$
@@
2 Initialize $\bm{K}_{0}\coloneqq(\bm{K}_{0}^{1},\ldots,\bm{K}_{0}^{N_{P}})$ and $k\coloneqq 0$;
@@ 8            $k\coloneqq k+1$;
The relationship between $BR$ and $BR_{\Phi}$ implies that a GFNE of $\mathcal{G}_{\infty}^{\text{LQ}}$ can be obtained analytically from a GNE of $\mathcal{G}_{\infty}^{\Phi}$. This allows us to adapt the BRD-GFNE procedure from Algorithm 3 into a procedure for GFNE seeking in constrained infinite-horizon dynamic games based on the mappings $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$. This system-level best-response dynamics (SLS-BRD) approach is given in Algorithm 4. We remark on some technical aspects:
  • The pair $(\bm{\Phi_{x,k}}^{p},\bm{\Phi_{u,k}}^{p})$ is the parametrisation of $p$-th player’s policy after $k\in\mathbb{N}$ updates, that is, the signals $\bm{\Phi_{x,k}}^{p}=(\Phi_{x,k,t}^{p})_{t\in\mathbb{N}}$ and $\bm{\Phi_{u,k}}^{p}=(\Phi_{u,k,t}^{p})_{t\in\mathbb{N}}$. The index $k\in\mathbb{N}$ should not be mistaken for the stage index $t\in\mathbb{N}$.
  • Updating the policy to $\bm{K}_{k+1}^{p}$ consists of employing the Corollary 1.1 with the updated maps $(\bm{\Phi_{x,k+1}}^{p},\bm{\Phi_{u,k+1}}^{p})$.
The SLS-BRD induces an operator $T_{\Phi}=(1-\eta)I+\eta BR_{\Phi}$, which defines the global update $\bm{\Phi_{u,k+1}}$ from $\bm{\Phi_{u,k}}$. As before, $T_{\Phi}$ and $BR_{\Phi}$ share the same fixed-points: The GNEs $\Omega_{\mathcal{G}_{\infty}^{\Phi}}$. From Theorem 3, convergence to a response $\bm{\Phi_{u}}^{\star}\in T_{\Phi}(\bm{\Phi_{u}}^{\star})$ then implies convergence to a policy $\bm{K}^{\star}\in\Omega_{\mathcal{G}_{\infty}}^{\bm{K}}$. Hence, this learning dynamics is a formal procedure for GFNE seeking.
@@
The convergence of the Algorithm 4 depends on $BR_{\Phi}$ (or $\widehat{BR}_{\Phi}$, for its tractable version) being at least nonexpansive (Lemmas 12). Formally,
Corollary 3.1.
Let $\widehat{BR}_{\Phi}:\bm{\mathcal{C}}_{u}\rightrightarrows\bm{\mathcal{C}}_{u}$ be $L_{\widehat{BR}_{\Phi}}$-Lipschitz, with $L_{\widehat{BR}_{\Phi}}<1$. Then, the SLS-BRD $\bm{\Phi_{u,k+1}}=\widehat{T}_{\Phi}(\bm{\Phi_{u,k}})$ converge to the unique $\varepsilon$-GNE $\bm{\Phi_{u}}^{\star}\in\Omega_{\mathcal{G}_{\infty}^{\Phi}}^{\varepsilon}$ with rate
$\cfrac{\|\bm{\Phi_{u,k}}-\bm{\Phi_{u}}^{\star}\|_{\ell_{2}}}{\|\bm{\Phi_{u,0}}-\bm{\Phi_{u}}^{\star}\|_{\ell_{2}}}\leq\big{(}(1{-}\eta)-\eta L_{\widehat{BR}_{\Phi}}\big{)}^{k}$ (40)
@@
$\underset{\bm{\Phi_{u}}^{p}\in U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})}{\text{minimize}}~{}$ $\mathrm{Tr}\big{[}{\bm{\Phi_{u}}^{p}}^{{\mkern-1.5mu\mathsf{T}}}\bm{H}^{pp}\bm{\Phi_{u}}^{p}$
$\quad+2(\textstyle\sum_{\tilde{p}\in\mathcal{P}\backslash\{p\}}\bm{H}^{p\tilde{p}}\bm{\Phi_{u}}^{\tilde{p}}+\bm{H}^{p0})^{{\mkern-1.5mu\mathsf{T}}}\bm{\Phi_{u}}^{p}\big{]}$ (42)
The maps $BR^{p}(\bm{\Phi_{u}}^{-p})$, $p\in\mathcal{P}$, thus correspond to the solution of quadratic programs with convex constraints $U_{\Phi}^{p}(\bm{\Phi_{u}}^{-p})$, for $\bm{\Phi_{u}}^{-p}\in\bm{\mathcal{C}}_{u}^{-p}$. In the case of $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|w_{t}\|_{\infty}\leq 1\}$ and no structural constraints ($\bm{S}_{x}=\bm{S}_{u}^{p}=I$), we have
@@
$\Phi_{x,{n+1}}=A\Phi_{x,n}{+}\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}\Phi^{\tilde{p}}_{u,n},~{}\Phi_{x,1}=I_{N_{x}},$
$\big{\|}([G_{x}]_{i}\bar{\Phi}_{x})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq 1,~{}i=1,\ldots,N_{\mathcal{X}};$
$\big{\|}([G_{u}^{p}]_{j}\bar{\Phi}_{u}^{p})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq 1,~{}j=1,\ldots,N_{\mathcal{U}^{p}};$
$\big{\|}([G_{\mathcal{G}}]_{l}\bar{\Phi}_{u})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}\leq 1,~{}l=1,\ldots,N_{\mathcal{U}_{\mathcal{G}}}\}.$
Using standard techniques from optimisation, the constraints Eq. (43) can be incorporated into Problem (42) as linear inequalities. The best-responses mappings $\{\widehat{BR}_{\Phi}^{p}\}_{p\in\mathcal{P}}$, are thus piecewise affine in $\bm{\Phi_{u}}^{-p}\in\bm{\mathcal{C}}_{u}^{-p}$ [45]. Consequently, also $\widehat{BR}_{\Phi}$ must be piecewise affine: A local Lipschitz constant can then be derived for each region of $\bm{\mathcal{C}}_{u}^{-p}$ that leads to a subset of the operational constraints being active. For non-generalised games, this fact can be exploited to derive a global Lipschitz constant for $\widehat{BR}_{\Phi}$.
@@
Consider $\mathcal{X}=\mathbb{R}^{N_{x}}$, $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$, and $\mathcal{U}^{p}=\{u^{p}_{t}\in\mathbb{R}^{N_{u}^{p}}:G_{u}^{p}u^{p}_{t}\preceq\mathbf{1}_{N_{\mathcal{U}}^{p}}\}$ with $G_{u}^{p}$ full-row-rank. Then, the best-response map $\widehat{BR}_{\Phi}$ is $L_{\widehat{BR}_{\Phi}}$-Lipschitz with
$L_{\widehat{BR}_{\Phi}}^{2}=\textstyle\sum_{p\in\mathcal{P}}(L_{\widehat{BR}_{\Phi}}^{p})^{2},$ (44)
given the player-specific $L_{\widehat{BR}_{\Phi}}^{p}=\sigma_{\max}(\bm{H}^{p,-p})/\sigma_{\min}(\bm{H}^{pp})$.
@@
The proof of Theorem 4 is extensive: The reader is referred to the Supplementary Material for the full details. The constants $L_{\widehat{BR}_{\Phi}}^{p}$ ($p\in\mathcal{P}$) quantify the ability of each player to optimize its objective in face of its rivals’ interference. Importantly, this Lipschitz constant is not tight and $L_{\widehat{BR}_{\Phi}}<1$ using Eq. (44) is only a sufficient (but not necessary) condition for Corollary 3.1 to hold in practice. Regardless, it highlights some intuitive, but non-trivial, facts about the convergence of the SLS-BRD:
@@
  • The condition $L_{\widehat{BR}_{\Phi}}<1$ is satisfied if the block-operator $\bm{H}=[H^{p\tilde{p}}]_{p,\tilde{p}\in\mathcal{P}}$ is diagonally dominant. This highlights the relationship between the SLS-BRD and classical Jacobi iterative methods, where diagonal dominance of the linear system being solved is a known requirement.
@@
  • The convergence rates of the SLS-BRD depend directly on $\|D^{pp}\|_{2}$ ($\forall p$) through $\sigma_{\min}(\bm{H}^{pp})=\sigma_{\min}^{2}(\bm{C}^{p}\bm{F_{\Phi}}^{p}+\bm{D}^{pp})$. As such, faster convergence is expected for games in which players apply strong penalties to their own actions.
  • The convergence rates of the SLS-BRD depend on the number of players since $L_{\widehat{BR}_{\Phi}}^{p}>0$ for all $p\in\mathcal{P}$. In large-scale games, players might need to apply strong penalties to their actions (through $D^{pp}$) to ensure convergence.